OpenStudy (anonymous):

tickets to a local movie were sold at 4.00$ for adults and 2.50$ for students. If 59 tickets were sold for a total of 222.50$, how many students were sold?

OpenStudy (lgbasallote):

"students were sold"<--funny anyway....let x be the number of students and y be the number of adults

OpenStudy (lgbasallote):

it says the total tickets was 59 so you can write that as x + y = 59 agree?

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

then, it says the total cost was 222.50 so... 2.50x + 4y = 222.50 because eachs tudent ticket is 2.5 worth and each adult ticket is 4 worth got it?

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

so now use those two equations x + y = 59 2.5x + 4y = 222.5 solve for x

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

@lgbasallote